Scores for a common standardized college aptitude test are normally distributed with a mean of 491 and a standard deviation of 102. Randomly selected men are given a Test Preparation Course before taking this test. Assume, for sake of argument, that the preparation course has no effect.

If 1 of the men is randomly selected, find the probability that his score is at least 555.6.
P(X > 555.6) =  
Enter your answer as a number accurate to 4 decimal places.

If 9 of the men are randomly selected, find the probability that their mean score is at least 555.6.
P(M > 555.6) =  
Enter your answer as a number accurate to 4 decimal places.

Assume that any probability less than 5% is sufficient evidence to conclude that the preparation course does help men do better. If the random sample of 9 men does result in a mean score of 555.6, is there strong evidence to support the claim that the course is actually effective?

  • Yes. The probability indicates that is is (highly ?) unlikely that by chance, a randomly selected group of students would get a mean as high as 555.6.
  • No. The probability indicates that is is possible by chance alone to randomly select a group of students with a mean as high as 555.6
Answer #1

Answer)

As the data is normally distributed we can use standard normal z table to estimate the answers

Z = (x-mean)/s.d

Given mean = 491

S.d = 102

A)

P(x<555.6)

Z = (555.6-491)/102 = 0.63

From z table, P(z<0.63) = 0.7357

B)

For a sample z = (x-mean)/(s.d/√n)

N =9

Z = (555.6-491)/(102/√9)

Z = 1.9

From z table, P(z<1.9) = 0.9713